OpenStudy (anonymous):

A ball is thrown upwards with an initial velocity of 10 m·s−1. Determine the maximum height reached above the thrower's hand. Determine the time it takes the ball to reach its maximum height.

OpenStudy (anonymous):

Is it thrown vertically upwards?

OpenStudy (anonymous):

If yes then, v^2 =u^2 +2as 0^2 = 10^2 + 2*(-9.81) *s

OpenStudy (anonymous):

Solve for s ... where s= max height reached by the ball.

OpenStudy (anonymous):

YA r8.......

OpenStudy (anonymous):

To calculate time u can use: s= (u+v)/2 *t

OpenStudy (anonymous):

Or, v=u+at